OpenStudy (anonymous):

Use "t results" to solve cosx-sinx=1 for 0<=x<=360

OpenStudy (zzr0ck3r):

what is "t results"?

OpenStudy (anonymous):

the t formulas. for example sinx = 2t/(1+t^2)

OpenStudy (zzr0ck3r):

what is the one for cos?

OpenStudy (anonymous):

(1-t^2)/(1+t^2)

OpenStudy (zzr0ck3r):

2t/(1+t^2)-(1-t^2)/(1+t^2) = 1 then?

OpenStudy (jhannybean):

I've never heard of these O_o

OpenStudy (anonymous):

lol i think t = tanx here

OpenStudy (zzr0ck3r):

no idea what this is...sorry

OpenStudy (anonymous):

t= tan(x)/(2)

OpenStudy (zzr0ck3r):

2t/(1+t^2)-(1-t^2)/(1+t^2) = 1 multiply by 1+t^2 2t-1+t^2 = 1+t^2 2t-1=1 2t=2 t=1 tan(x)/2=1 x=arctan(2)?

OpenStudy (anonymous):

it gives t = 1 so tan(x/2) = 1 x/2 = pi/4, 5pi/4 x = pi/2, 5pi/2

OpenStudy (anonymous):

It's the half angle (x/2) @zzr0ck3r

OpenStudy (zzr0ck3r):

ahh its tan(x/2)?

OpenStudy (zzr0ck3r):

ahh

OpenStudy (anonymous):

gotcha ? @hothand621

OpenStudy (zzr0ck3r):

2t/(1+t^2)-(1-t^2)/(1+t^2) = 1 multiply by 1+t^2 2t-1+t^2 = 1+t^2 2t-1=1 2t=2 t=1 tan(x/2)=1 x/2=arctan(1) find the values for arctan(1) which are pi/4 and 5pi/4, so x/2 = pi/4 so x = pi/2 and x/2 = 5pi/4 so x = 5pi/2

OpenStudy (zzr0ck3r):

had to finish what I started:)